How would i calculate the force required to balance a beam, if on the other side there are two forces acting upon it and different distances, with one force being upwards and other downwards?

For example on the left side there is a downwards force1 of 6N and 0.2m away from the pivot, and an upwards force of 10N and 0.1m away from the pivot. On the other side there is an unknown force 0.4m away from the pivot- what is the force required to balance the beam?
1
Expert's answer
2018-01-26T09:43:07-0500
When you need to find one force (other are known) you can use the fact that the sum of forces is zero.
In our case:
Let upwards be the positive direction. Thus,
-F_1+F_2+F=0
F=F_1-F_2=6-10=-4 N.
Thus, the force required to balance the beam is 4 N downwards.